# Optimal control transport of neutral atoms in optical tweezers at finite temperature

### v1

@@ Abstract
The transport of neutral atoms in Rydberg quantum computers is a crucial step of the initial arrangement of the grid as well as to the dynamic connectivity, recently successfully demonstrated [1]. We study the application of optimal control and the quantum speed limit for the transport of neutral atoms in optical tweezers at finite temperatures and analyze how laser noise affects transport fidelity. Open-loop optimal control significantly enhances transport fidelity, achieving an improvement up to 89% for the lowest analyzed temperature of $1\,\mu$K for a distance of three micrometers. Furthermore, we simulate how the transport fidelity behaves in release-and-capture measurements, which are realizable in the experiment to estimate transport efficiency and implement closed-loop optimal control.
@@
Neutral atom quantum processing units based on optical tweezers [2, 3, 4, 5, 6, 7, 8] open the possibility to modify the connectivity of qubits at runtime of the algorithm. In the pioneering experiments in Ref. [9, 1], coherent transport of atoms through the movement of optical tweezers has been demonstrated. Ensuring high-fidelity atom movement is crucial not only for enabling non-local connectivity but also for the necessary reshuffling of atoms in the initial array arrangement, after the stochastic loading of atoms within the optical tweezers [10, 11, 12, 13, 14]. A fast high-fidelity transport can be achieved via quantum optimal control (QOC) [15, 16, 17, 18]. Optimal control has emerged as an essential tool for optimizing quantum gates as well [19, 20, 21, 1]. QOC allows for either open-loop optimization using numerical simulations or closed-loop optimization directly integrated with the experimental setup. A variety of optimal control methods is available via open-source package [22] as the dressed chopped random basis (dCRAB) algorithm [23, 24, 25].
[…]
@@
Figure 1: Setup for atoms transport. (a) An array of neutral atoms after the initial stochastic loading. Empty traps are switched off and atom transport along one row of the array is possible via optical tweezer generated by acousto-optical deflectors (AODs). (b-c) We aim to move the atom at position $r_{0}=0\mu\mathrm{m}$ to the position $r_{\mathrm{f}}=3\mu\mathrm{m}$, a vacant spot in the example. (d) We perform open-loop optimal control via the time-averaged figure of merit $J_{\mathrm{avg}}$; closed-loop optimal control is anticipated by taking into account additional experimental constraints in our simulation, especially on the pulse shape $r(t)$.
@@
We design an open-loop optimal transport of neutral atoms via optical tweezer at finite temperature to closely replicate experimental conditions: indeed, in typical experiments, atoms are trapped in Gaussian-shape potential formed by optical tweezers and each atom can be represented as a thermal state at a temperature approximately around $10\mu\mathrm{K}$ [27]. After the stochastic loading of the atoms, empty traps are switched off, as depicted in Fig. 1(a). We have control to move tweezers within a given row through acousto-optical deflectors (AODs). We reshuffle the atoms to occupy the vacant positions, thereby achieving the desired layout. The movement must be executed at sufficient fidelity and speed to prevent loss and avoid any temperature alterations [28]. We take into consideration a variety of experimental constraints for optical tweezers implemented via acousto-optical deflectors, e.g., a finite frequency bandwidth, and piecewise quadratic pulse shapes, both motivated by the experiment in Ref. [27, 29, 30]. We also introduce noise into the modeling of the tweezer trap, i.e. for its depth, waist, and position. We find that, with our setup, the quantum speed limit for moving a Strontium-88 atom over a distance of three micrometers is around $10\,\mu$s. Furthermore, at the lowest analyzed temperature, a notable enhancement in the figure of merit is observed, reaching a factor of 89%. We examine the impact of laser noise on the trap on the figure of merit. The release-and-capture protocol is the standard approach to verify the quality of the atom cooling [27], which also applies to the evaluation of the heating during transport. We predict which level of precision of the figure of merit can be observed experimentally by running the release-and-capture simulations after the optimal pulses.
@@ I Optimal control setup
$\rho(T)$ $=\sum_{i=1}^{n_{\rm\scriptscriptstyle c}}p_{i}\ket{\psi_{i}}\bra{\psi_{i}}\,,$ (1)
where $p_{i}=p_{i}(T)=\mathrm{e}^{-E_{i}/k_{b}T}$ are the Boltzmann weights at temperature $T$. We introduce a cutoff in the number of states $n_{\rm\scriptscriptstyle c}$ to represent the finite temperature density matrix. The convergence of the results with the cutoff $n_{\rm\scriptscriptstyle c}$ is discussed in appendix  A. We model the optical tweezer traps as Gaussian potentials of the following form:
$V(t)=U_{0}\mathrm{e}^{\frac{-2r(t)^{2}}{w_{0}^{2}}}\,,$ (2)
where $U_{0}$ is the trap depth, $w_{0}$ the beam waist and $r$ the time-dependent center of the trap [11, 6, 29]. In the experiment, the pulse $r(t)$ tuned to move the atom is generated via arbitrary waveform generators (AWG), which imposes constraints on the available pulse shapes, such as the piecewise quadratic shape [31]. Additionally, we consider laser noise effects and finite bandwidth during the optimization process. We describe the numerical simulation for computing the time-evolution of the thermal state under time-dependent pulses in appendix A and integrate quantum optimal control using the open-source suite QuOCS [22]. We employ the dCRAB algorithm [25, 24]: the key idea is to minimize a figure of merit, i.e., $J(r)$, by expanding the control pulse $r(t)$ as a sum of truncated randomized basis $r_{i}(t)$. In more detail, for a super-iteration $j$, the control pulse is constructed by finding the optimal set of coefficients $c_{i}$:
@@
A common selection for basis decomposition is the use of trigonometric functions, specifically the Fourier basis; however, the choice of bases can vary depending on the nature of the control problem, with certain bases proving more effective than others [32]. Specifically, we employed the sinc basis for optimizing the pulse $r(t)$. The figure of merit is the infidelity between the density matrices at the end of the pulse $\rho_{\mathrm{f}}(r)$ and the density matrix of the initial finite-temperature state shifted to the final position $\rho_{0}(r_{\mathrm{f}})$, i.e.,
@@
In the following analysis, we target Strontium-88 atoms which have a mass of $m=1.46\times 10^{-25}\,\text{kg}$ and we fix a distance between the traps in the array of $d=3\mu\text{m}$. For the Gaussian-shaped trap, we assume a trap depth of $U_{0}=-1\text{mK}$ with a beam waist $w_{0}=0.5\mu\text{m}$. The coordinate in position-space is set via the frequency of the AOD, where the conversion is $13\mu\text{m}=$ 1MHz. The noise for the position, waist, and depth of the Gaussian potential is modeled as follows: the amplitude of the laser controlling the trap depth exhibits a noise $\delta S$ with a spectrum following a power-law distribution given by $S(f)=A_{L}/f^{1/2}$, i.e., the relative intensity noise (RIN). We know from measurement [33] that the amplitude is about $A_{L}=10^{-11}$. The noise of the laser itself is not significant for the transport, but the beam is typically modified by fiber couplings, the optical system, and especially the AOD; the AOD with its raise time of $6\mu$s modifies the beam. This is considered as additional noise $\delta U_{0}$ in the trap depth, introduced by random sinusoidal fluctuations with an amplitude of 1% and random frequencies below 100100100\,100kHz, as well as random sinusoidal fluctuations with an amplitude of 5% and random frequencies exceeding 100100100\,100kHz. For the trap waist, we assume sinusoidal noise $\delta\omega_{0}$ with a frequency of $1/6\,\mu$s motivated by the raise time of the AOD and an amplitude of 1%. Regarding the trap position, we assume sinusoidal noise $\delta r$ with an amplitude of $0.01\,\mu$m and random frequencies ranging between $50-150\,$MHz.
@@ II Time-optimal pulse for atom transport
where $t_{\mathrm{p}}$ is the duration of the pulse. Since the figure of merit defined in Eq. (4) oscillates at the end of the transport, we evolve the system for an additional fixed $t_{\mathrm{c}}=10\,\mu\text{s}$ to compute the time-average $J_{\mathrm{avg}}$. Thus, the evolution time is $t_{\mathrm{t}}=t_{\mathrm{p}}+t_{\mathrm{c}}$. The optimization workflow is summarized in Fig. 1(d), with an example of a piecewise quadratic pulse represented by the black line. We examine the time-average of the figure of merit, $J_{\mathrm{avg}}$, as a function of the pulse duration $t_{\mathrm{p}}$ for the piecewise quadratic pulse, considering various temperatures for the thermal state in Eq. (1), specifically for $T=1\,\mu\text{K},10\,\mu\text{K},30\,\mu$K. The results are shown in Fig. 2(a). The eigenstates of the Gaussian potential lead to typical frequencies and timescales in the system, which is shown as periodic minima in the figure of merit; this result agrees with the observation for transport in optical lattices, see Ref. [34]. For the strontium-88 atom and for the selected trap depth, the periodicity is given by $2\tau\approx 10\,\mu$s with
@@
where $E_{1}$ and $E_{0}$ are the eigenenergies of the first excited and ground state of the Gaussian trap, respectively. Using the piecewise quadratic pulse, we achieve a figure of merit $J_{\mathrm{avg}}<10^{-2}$ for $t_{\mathrm{p}}>20\,\mu$s, for the considered temperatures as shown in Fig. 2(a).
[…]
@@
Figure 2: Figure of merit as a function of transport time for different temperatures. (a) For a fixed pulse duration $t_{\mathrm{p}}$, the center of the Gaussian-shaped trap is moved via a piecewise quadratic pulse $r_{{\mathrm{pq}}}$ for the distance of $3\mu$m. We observe periodic minima with a frequency of $2\tau=10\,\mu$s. When these minima occur, the pulse length $t_{\mathrm{p}}$ is precisely tuned to enable the wavefunction to complete its oscillation period. (b) The pulse $r_{{\mathrm{oc}}}$ is obtained with optimal control with $r_{{\mathrm{pq}}}$ as initial guess. The optimal control removes the dependency on the wavefunction oscillations and reaches a figure of merit $J_{\mathrm{avg}}$ which monotonically decreases with the pulse time $t_{p}$. We observe the quantum speed limit around $2\tau$. The figure of merit depends on the temperature $T$, i.e., higher temperatures result in a worse $J_{\mathrm{avg}}$ for the optimal pulse.
@@
In Fig. 2(b) we now optimize the pulse shape via QOC: we analyze $J_{\mathrm{avg}}$ as a function of the optimal pulse with duration $t_{\mathrm{p}}$. The initial guess for the optimization is the piecewise quadratic pulse $r_{\mathrm{pq}}$. This data shows the typical feature of the quantum speed limit in an optimal control problem [35, 34]: we extract the quantum speed limit from Fig. 2(b) as $2\tau$ where the figure of merit does not improve anymore. This estimate is in good agreement with the timescale of the gap in Eq. (6). The optimal control optimization eliminates the figure of merit oscillations that were present in the piecewise quadratic case and results in $J_{\mathrm{avg}}=10^{-2}$ at around $t_{\mathrm{p}}=10-12\,\mu$s depending on the temperature $T$. We observe that the temperature significantly influences the figure of merit for atom transport. Higher temperatures lead to a worse figure of merit for the optimal pulse, with no room for further optimization. The relative improvement of the optimal control solution for the piecewise quadratic pulse is illustrated in Fig. 3 as a function of the pulse duration. The improvement holds particular significance at pulse duration $t_{\mathrm{p}}$ where minima were evident for $r_{\mathrm{pq}}$ and as close as possible to the quantum speed limit. The optimal control can explore interesting solutions in such scenarios [15]. Specifically at $t_{\mathrm{p}}=11\,\mu$s, the figure of merit can be improved by a factor of $89\%,58\%,66\%$ for $T=1\,\mu\text{K},10\,\mu\text{K},30\,\mu$K, respectively. Furthermore, we verify that a comparable improvement is observed when we impose the experimental constraint on the pulse shape fitting it with piecewise quadratic segments, each spanning $10\,\mu$s.
[…]

### v2

@@ Abstract
The transport of neutral atoms in Rydberg quantum computers is a crucial step of the initial arrangement of the grid as well as to the dynamic connectivity, recently successfully demonstrated.We study the application of optimal control and the quantum speed limit for the transport of neutral atoms in optical tweezers at finite temperatures and analyze how laser noise affects transport fidelity. Open-loop optimal control significantly enhances transport fidelity, achieving an improvement up to 89% for the lowest analyzed temperature of $1\,\mu$K for a distance of three micrometers. Furthermore, we simulate how the transport fidelity behaves in release-and-capture measurements, which are realizable in the experiment to estimate transport efficiency and implement closed-loop optimal control.
@@
Neutral atom quantum processing units based on optical tweezers [1, 2, 3, 4, 5, 6, 7] open the possibility to modify the connectivity of qubits at runtime of the algorithm. In the pioneering experiments in Ref. [8, 9], coherent transport of atoms through the movement of optical tweezers has been demonstrated. Ensuring high-fidelity atom movement is crucial not only for enabling non-local connectivity but also for the necessary reshuffling of atoms in the initial array arrangement, after the stochastic loading of atoms within the optical tweezers [10, 11, 12, 13, 14, 15]. A fast high-fidelity transport can be achieved via quantum optimal control (QOC) [16, 17, 18, 19]. Optimal control has emerged as an essential tool for optimizing quantum gates as well [20, 21, 22, 9]. QOC allows for either open-loop optimization using numerical simulations or closed-loop optimization directly integrated with the experimental setup. A variety of optimal control methods is available via open-source package [23] as the dressed chopped random basis (dCRAB) algorithm [24, 25, 26].
@@
Figure 1: Setup for atoms transport. (a) An array of neutral atoms after the initial stochastic loading. Empty traps are switched off and atom transport along one row of the array is possible via optical tweezer generated by acousto-optical deflectors (AODs). (b-c) We aim to move the atom at position $r_{0}=0\mu\mathrm{m}$ to the position $r_{\mathrm{f}}=3\mu\mathrm{m}$, a vacant spot in the example. (d) We perform open-loop optimal control via the time-averaged figure of merit $J_{\mathrm{avg}}$; closed-loop optimal control is anticipated by taking into account additional experimental constraints in our simulation, especially on the pulse shape $r(t)$ that controls the position of the trap.
@@
We design an open-loop optimal transport of neutral atoms via optical tweezer at finite temperature to closely replicate experimental conditions: indeed, in typical experiments, atoms are trapped in Gaussian-shape potential formed by optical tweezers and each atom can be represented as a thermal state at a temperature approximately around $10\mu\mathrm{K}$ [27]. After the stochastic loading of the atoms, empty traps are switched off, as depicted in Fig. 1(a). We have control to move tweezers within a given row through acousto-optical deflectors (AODs). We reshuffle the atoms to occupy the vacant positions, thereby achieving the desired layout. The movement must be executed at sufficient fidelity and speed to prevent loss and avoid any temperature alterations [28]. We take into consideration a variety of experimental constraints for optical tweezers implemented via acousto-optical deflectors, e.g., a finite frequency bandwidth, and piecewise quadratic pulse shapes, both motivated by the experiment in Ref. [27, 29, 30]. We also introduce noise into the modeling of the tweezer trap, i.e. for its depth, waist, and position. We find that, with our setup, the quantum speed limit for moving a Strontium-88 atom over a distance of three micrometers is around $10\,\mu$s. The quantum speed limit relates to the minimum time required to accomplish a physical process. For transport, we extrapolate this as the time when it is no longer possible to improve the figure of merit, in agreement with [31]. Furthermore, at the lowest analyzed temperature, a notable enhancement in the figure of merit is observed, reaching a factor of 89%. We examine the impact of laser noise on the trap on the figure of merit. The release-and-capture protocol is the standard approach to verify the quality of the atom cooling [27], which also applies to the evaluation of the heating during transport. We predict which level of precision of the figure of merit can be observed experimentally by running the release-and-capture simulations after the optimal pulses. The proposed analysis applies only to the initialisation of the array. Extending it to mid-circuit movement of atoms would require encoding the internal degrees of freedom of the atoms and decoherence effects, which is beyond the scope of this work.
@@ I Optimal control setup
$\rho(T)$ $=\sum_{i=1}^{N_{s}}p_{i}\ket{\psi_{i}}\bra{\psi_{i}}\,,$ (1)
where $p_{i}=p_{i}(T)=\mathrm{e}^{-E_{i}/k_{b}T}$ are the Boltzmann weights at temperature $T$ with $E_{i}$ the energy of the $i$-state. We introduce a cutoff in the number of states $N_{s}$ to represent the finite temperature density matrix. The convergence of the results with the cutoff $N_{s}$ is discussed in appendix  A. We model the optical tweezer traps as Gaussian potentials of the following form:
$V(t)=U_{0}\mathrm{e}^{\frac{-2(x-r(t))^{2}}{w_{0}^{2}}}\,,$ (2)
where $U_{0}$ is the trap depth, $w_{0}$ the beam waist and $r$ the time-dependent center of the trap [12, 5, 29]. In the experiment, the pulse $r(t)$ tuned to move the atom is generated via arbitrary waveform generators (AWG), which imposes constraints on the available pulse shapes, such as the piecewise quadratic shape [32]. Additionally, we consider laser noise effects and finite bandwidth during the optimization process. We describe the numerical simulation for computing the time-evolution of the thermal state under time-dependent pulses in appendix A and integrate quantum optimal control using the open-source suite QuOCS [23]. We employ the dCRAB algorithm [26, 25]: the key idea is to construct a pulse $r(t)$ as a sum of randomly assembled basis elements $r_{i}(t)$ to minimize a given figure of merit $J$. The optimization involves different superiterations, each simultaneously optimizing $N_{c}$ coefficients of the basis. The full pulse at the superiteration $j$ can be written as
@@
where $r^{j-1}$ is the converged solution of the previous superiteration and $\{c_{i}\}_{i=0}^{N_{c}}$ is the set of coefficients to optimize. A common selection for basis decomposition is the use of trigonometric functions, specifically the Fourier basis; however, the choice of bases can vary depending on the nature of the control problem, with certain bases proving more effective than others [33]. Specifically, we employed the sinc basis for optimizing the pulse $r(t)$. The figure of merit is the infidelity between the density matrices at the end of the pulse $\rho_{\mathrm{f}}(r)$ and the density matrix of the initial finite-temperature state shifted to the final position $\rho_{0}(r_{\mathrm{f}})$, i.e.,
@@
This metric is a reliable distance measure between any two states [34, 35]. A worse figure of merit indicates that at the end of the transport, the atom has some residual kinetic energy, which directly relates to an increase in the atom’s temperature. This statement holds as the atoms starts in a thermal state and its temperature can only increase during a unitary evolution.
In the following analysis, we target Strontium-88 atoms which have a mass of $m=1.46\times 10^{-25}\,\text{kg}$ and we fix a distance between the traps in the array of $d=3\mu\text{m}$. The analysis for different distances is discussed in Appendix  B. For the Gaussian-shaped trap, we assume a trap depth of $U_{0}=-1\text{mK}$ with a beam waist $w_{0}=0.5\mu\text{m}$. The coordinate in position-space is set via the frequency of the AOD, where the conversion is $13\mu\text{m}=$ 1MHz. The noise for the position, waist, and depth of the Gaussian potential is modeled as follows: the amplitude of the laser controlling the trap depth exhibits a noise $\delta S$ with a spectrum following a power-law distribution given by $S(f)=A_{L}/f^{1/2}$, i.e., the relative intensity noise (RIN). We know from measurement [36] that the amplitude is about $A_{L}=10^{-11}$. The noise of the laser itself is not significant for the transport, but the beam is typically modified by fiber couplings, the optical system, and especially the AOD; the AOD with its raise time of $6\mu$s modifies the beam. This is considered as additional noise $\delta U_{0}$ in the trap depth, introduced by random sinusoidal fluctuations with an amplitude of 1% and random frequencies below 100100100\,100kHz, as well as random sinusoidal fluctuations with an amplitude of 5% and random frequencies exceeding 100100100\,100kHz. For the trap waist, we assume sinusoidal noise $\delta\omega_{0}$ with a frequency of $1/6\,\mu$s motivated by the raise time of the AOD and an amplitude of 1%. Regarding the trap position, we assume sinusoidal noise $\delta r$ with an amplitude of $0.01\,\mu$m and random frequencies ranging between $50-150\,$MHz.
@@ II Time-optimal pulse for atom transport
where $t_{\mathrm{p}}$ is the duration of the pulse. Since the figure of merit defined in Eq. (4) oscillates at the end of the transport, we evolve the system for an additional fixed $t_{\mathrm{c}}=10\,\mu\text{s}$ to compute the time-average $J_{\mathrm{avg}}$. Thus, the evolution time is $t_{\mathrm{t}}=t_{\mathrm{p}}+t_{\mathrm{c}}$. The optimization workflow is summarized in Fig. 1(d), with an example of a piecewise quadratic pulse represented by the black line. We examine the time-average of the figure of merit, $J_{\mathrm{avg}}$, as a function of the pulse duration $t_{\mathrm{p}}$ for the piecewise quadratic pulse, considering various temperatures for the thermal state in Eq. (1), specifically for $T=1\,\mu\text{K},10\,\mu\text{K},30\,\mu$K. The results are shown in Fig. 2(a). The eigenstates of the Gaussian potential lead to typical frequencies and timescales in the system, which is shown as periodic minima in the figure of merit; this result agrees with the observation for transport in optical lattices, see Ref. [31]. For the strontium-88 atom and for the selected trap parameters, the periodicity is given by $2\tau\approx 10\,\mu$s with
@@
where $E_{1}$ and $E_{0}$ are the eigenenergies of the first excited and ground state of the Gaussian trap, respectively. Using the piecewise quadratic pulse, we achieve a figure of merit $J_{\mathrm{avg}}<10^{-2}$ for $t_{\mathrm{p}}\geq 20\,\mu$s, for the considered temperatures as shown at the top of Fig. 2(a). We define the first transport time that satisfies this condition on $J_{\mathrm{avg}}$ as $t_{\text{min}}$. For $T=1\,\mu$K, we have $t_{\text{min}}^{\text{pq}}=20\,\mu$s.
@@
Figure 2: Figure of merit as a function of transport time for different temperatures. (a) On top, for a fixed pulse duration $t_{\mathrm{p}}$, the center of the Gaussian-shaped trap is moved via a piecewise quadratic pulse $r_{{\mathrm{pq}}}$ for the distance of $3\mu$m. We observe periodic minima with a frequency of $2\tau=10\,\mu$s. When these minima occur, the pulse length $t_{\mathrm{p}}$ is precisely tuned to enable the wavefunction to complete its oscillation period. At the bottom, the pulse $r_{{\mathrm{oc}}}$ is obtained with optimal control with $r_{{\mathrm{pq}}}$ as initial guess. The optimal control removes the dependency on the wavefunction oscillations and reaches a figure of merit $J_{\mathrm{avg}}$ which monotonically decreases with the pulse time $t_{p}$. We observe the quantum speed limit around $2\tau$. The figure of merit depends on the temperature $T$, i.e., higher temperatures result in a worse $J_{\mathrm{avg}}$ for the optimal pulse. (b) Difference between the optimized pulse $r_{\text{oc}}(t)$ and the initial guess $r_{\text{pq}}(t)$ for $t_{\text{p}}=11\,\mu$s at $T=1\,\mu$K. (c) The difference is bigger for $t_{\text{p}}=9\,\mu$s at the same temperature.
@@
At the bottom of Fig. 2(a) we show the results of the optimization of the pulse shape via QOC: we analyze $J_{\mathrm{avg}}$ as a function of the optimal pulse with duration $t_{\mathrm{p}}$. The initial guess for the optimization is the piecewise quadratic pulse $r_{\mathrm{pq}}$. This data shows the typical feature of the quantum speed limit in an optimal control problem [37, 31]: we extract the quantum speed limit from the bottom plot of Fig. 2(a) as $2\tau$ where the figure of merit does not improve anymore. This estimate is in good agreement with the timescale of the gap in Eq. (6). From now on, we interchangeably refer to the quantum speed limit or $t_{\text{min}}$, the minimum time where the condition $J_{\mathrm{avg}}<10^{-2}$ holds. With optimal control and for $T=1\,\mu$K, we have $t_{\text{min}}^{\text{oc}}=11\,\mu$s. We note that this estimate is almost half the one with the piecewise pulse. In Fig. 2(b), we show the difference of an optimized pulse $r_{\text{oc}}$ with $t_{\text{p}}=11\,\mu$s with its initial guess. In this case, the optimal control solution adds a small correction to the piecewise quadratic shape to match the wavefunction’s periodicity. Another example is shown in Fig. 2(c). In general, the optimal control optimization eliminates the figure of merit oscillations that were present in the piecewise quadratic case and results in $J_{\mathrm{avg}}=10^{-2}$ at around $t_{\mathrm{p}}=10-12\,\mu$s depending on the temperature $T$. We observe that the temperature significantly influences the figure of merit for atom transport. Higher temperatures of the initial state lead to a worse figure of merit for the optimal pulse, with no room for further optimization. Moroever, we emphasize that the quantum speed limit depends on the fixed trapping parameters and we verify that it increases for a shallow trap. For instance, if we fix as trap depth $U_{0}=-0.5\,$mK, the quantum speed limit is $t_{\text{min}}^{\text{oc}}=14\,\mu$s. Surprisingly, for deeper traps the minimum time does not decrease but it converges to $t_{\text{min}}^{\text{oc}}=11\,\mu$s. The relative improvement of the optimal control solution for the piecewise quadratic pulse is illustrated in Fig. 3 as a function of the pulse duration. The improvement holds particular significance at pulse duration $t_{\mathrm{p}}$ where minima were evident for $r_{\mathrm{pq}}$ and as close as possible to the quantum speed limit. Specifically at $t_{\text{min}}^{\text{oc}}=11\,\mu$s, the figure of merit can be improved by a factor of $89\%,58\%,66\%$ for $T=1\,\mu\text{K},10\,\mu\text{K},30\,\mu$K, respectively. Furthermore, we verify that a comparable improvement is observed when we impose the experimental constraint on the pulse shape fitting it with piecewise quadratic segments, each spanning $10\,\mu$s.